Expand f(x) = "e^{2x} + 3x^2" in a Taylor series at a = 3 to the third derivative degree


1
Expert's answer
2021-10-15T09:49:02-0400

Solution:

"f(x)=e^{2x}+3x^2
\\f'(x)=2e^{2x}+6x
\\f''(x)=4e^{2x}+6
\\f'''(x)=8e^{2x}"

Then, at "x=3",

"f(3)=e^6+27
\\f'(3)=2e^6+18
\\f''(3)=4e^6+6
\\f'''(3)=8e^6"

We know that, Taylor series expansion at a = 3 to the third derivative degree is

"f(x)=P_3(x)=f(3)+f'(3)(x-3)+\dfrac{f''(3)}{2}(x-3)^2+\dfrac{f'''(3)}{3\times2}(x-3)^3"

"=e^6+27+(2e^6+18)(x-3)+\dfrac{4e^6+6}2(x-3)^2+\dfrac{8e^6}6(x-3)^3"

"=\dfrac{4e^6x^3-30e^6x^2+9x^2+78e^6x-69e^6}{3}"